A final year student after being interviewed at two companies, he assesses that his probability of getting an offer from company A is 0.8 and the probability that he gets offer from company B is 0.6. If, on the other hand he believes that the probability that he will get offers from both companies is 0.5. Obtain the probability that he will get
(i). at least one offer from these two companies
(ii). offer from neither company
(iii). offer from company A only 

Answer by ikleyn(52780)

.
A final year student after being interviewed at two companies, he assesses that his probability of getting an offer
from company A is 0.8 and the probability that he gets offer from company B is 0.6. If, on the other hand he believes
that the probability that he will get offers from both companies is 0.5. Obtain the probability that he will get
(i)    at least one offer from these two companies
(ii)   offer from neither company
(iii)  offer from company A only 
~~~~~~~~~~~~~~~~~~


Solution


(i)   P = P(A) + P(B) - P(A and B) = 0.8 + 0.6 - 0.5 = 0.9.



(ii)  In this case, the probability is the COMPLEMENT to the probability which is just found  in (i)

       P' = 1 - P = 1 - 0.9 = 0.1.



(iii)  P = P(A) - P(A and B) = 0.8 - 0.5 = 0.3.


Solved.     //     All questions are answered.